# Learning incomplete factorization preconditioners for GMRES

### v1

@@ Abstract
In this paper, we develop a data-driven approach to generate incomplete LU factorizations of large-scale sparse matrices. The learned approximate factorization is utilized as a preconditioner for the corresponding linear equation system in the GMRES method. Incomplete factorization methods are one of the most commonly applied algebraic preconditioners for sparse linear equation systems and are able to speed up the convergence of Krylov subspace methods. However, they are sensitive to hyper-parameters and might suffer from numerical breakdown or lead to slow convergence when not properly applied. We replace the typically hand-engineered algorithms with a graph neural network-based approach that is trained against data to predict an approximate factorization. This allows us to learn preconditioners tailored for a specific problem distribution. We analyze and empirically evaluate different loss functions to train the learned preconditioners and show their effectiveness to decrease the number of GMRES iterations and improve the spectral properties on our synthetic dataset. The code is available at https://github.com/paulhausner/neural-incomplete-factorization.
@@ 1 Introduction
The GMRES algorithm is one of the most popular iterative methods to solve large-scale and sparse linear equation systems of the form ${\bm{A}}{\bm{x}}={\bm{b}}$. Throughout the paper, we assume the system matrix ${\bm{A}}$ to be real-valued and full rank. Therefore, the unique solution ${\bm{A}}^{-1}{\bm{b}}$ exists. However, inverting the matrix directly, or equivalently computing a full matrix factorization, scales computationally poorly and suffers from numerical instabilities. Instead, iterative Krylov subspace methods (such as GMRES) which refine an approximation of the solution in each step are the most common approach to solving problems of this form. The convergence of these methods highly depends on the conditioning and singular value distribution of the system matrix. Therefore, the choice of appropriate preconditioner – designed to improve the system properties – is critical to obtain a fast and accurate solution for the equation system [10saad].
@@
One of the most common choices of preconditioners is the incomplete LU factorization (ILU). As the name suggests, the matrix ${\bm{A}}$ is factorized but not all elements of the factorization are computed to decrease the storage and computational time (benzi2002preconditioning; scott2023algorithms). In this work, we replace the classical computation of the LU factors with a graph neural network (GNN). We train the GNN to predict the corresponding incomplete factors of the matrix directly by training against data. The main contributions of our paper are the following:
  • We design a GNN model and train it to output a non-singular incomplete LU factorization.
@@
  • We theoretically analyze existing loss functions from the literature and showcase their connection to large and small singular values.
@@
  • Based on these insights, we derive a novel loss function to train learned preconditioners that combines the benefits of previous approaches.
@@
In numerical experiments, we showcase the effectiveness of our model to reduce GMRES iterations in combination with the different loss functions and validate the theoretical results on a synthetic dataset.
@@
Related work  Most similar to our work, hausner2023neural and li2023learning both learn an incomplete factorization for the conjugate gradient method. trifonov2024learning extend this to combine data-driven and classical algorithms by correcting the output of the incomplete Cholesky factorization with a learned component. However, all of these methods assume the matrix to be symmetric and positive definite and utilize the incomplete Cholesky factorization as the underlying preconditioning technique. In contrast, rather focus on the more general incomplete LU factorization, which does not require positive definiteness or symmetry, within the GMRES algorithm.
@@
For the class of sparse approximate inverse preconditioner, baankestad2024ising learn the sparsity pattern of the approximate inverse matrix. li2024generative instead propose to learn auto-encoder-based generative models to generate preconditioners. In contrast, this paper only focuses on learning factorized preconditioners that allow easy inversion.
@@ 2 Background
We first briefly introduce the GMRES algorithm and focus in particular on the design choice of preconditioners. Then, we describe graph neural networks which we use to parameterize the mapping to learn data-driven preconditioners.
@@ 2.1 GMRES algorithm
We focus on the GMRES algorithm (saad86gmres), a popular and widely adopted iterative Krylov subspace method to solve general linear equation systems of the form ${\bm{A}}{\bm{x}}={\bm{b}}$ where ${\bm{A}}$ is a $n\times n$ non-singular matrix denoted by the set GL$(n,\mathbb{R})$. The core idea of Krylov subspace methods is to iteratively refine the solution to the problem starting from an initial guess ${\bm{x}}_{0}$. At each iteration, the approximate solution is updated by minimizing the Euclidean norm of the residual within the corresponding Krylov subspace:
@@
where the $k$-th Krylov subspace $\mathcal{K}_{k}({\bm{A}},{\bm{r}}_{0})=\text{span}\{{\bm{r}}_{0},{\bm{A}}{\bm{r}}_{0},{\bm{A}}^{2}{\bm{r}}_{0}\dots,{\bm{A}}^{k-1}{\bm{r}}_{0}\}$ of ${\bm{A}}$ is generated by the initial residual ${\bm{r}}_{0}={\bm{b}}-{\bm{A}}{\bm{x}}_{0}$ (10saad). The new approximation ${\bm{x}}_{k}$ builds on the previous solution ${\bm{x}}_{k-1}$ by efficiently computing it within a progressively larger Krylov subspace by constructing the solution in a progressively larger subspace that captures more of the action of the matrix ${\bm{A}}$ on the residual. This expanding subspace offers more directions to approximate the solution, improving the conditioning and accuracy of the minimization problem over iterations. The efficiency arises because ${\bm{x}}_{k}$ is found by solving a smaller, reduced system in the subspace rather than the original larger system. The GMRES algorithm is guaranteed to find the exact solution to the linear equation system in at most $n$ steps.
@@
Each iteration consists of two main steps. In the first step of the algorithm, an orthonormal basis for the subspace $\mathcal{K}_{k}$ is computed. This can be achieved, for example, using Arnoldi’s algorithm which is based on the Gram-Schmidt procedure. In the second step, an unconstrained least squares problem is solved via the QR factorization of the subspace basis obtained in the first step. Based on this, the solution of Equation (1) can be constructed efficiently. The detailed algorithm is shown in Appendix A.
@@
Preconditioning  Arguably, the most important design choice for any Krylov subspace method is the choice of preconditioner. The goal of preconditioning is to replace the original linear equation system with a new preconditioned system that exhibits a better clustering of singular values, which in turn leads to faster convergence of the iterative scheme (golub2013matrix). For a non-singular and easy-to-invert matrix ${\bm{P}}$, the linear equation system ${\bm{A}}{\bm{P}}^{-1}{\bm{y}}={\bm{b}}$ is solved instead of the original problem. The solution to the original problem is then given by ${\bm{x}}={\bm{P}}^{-1}{\bm{y}}$ [^1] Here, right preconditioning is used, but other formulations use left- or split-preconditioning instead often leading to similar but not equivalent results. Constructing a preconditioner involves trading off the time required to compute the preconditioner itself and the resulting speed up in the iterative scheme (benzi2002preconditioning). Since the original system ${\bm{A}}$ is typically sparse, we restrict the preconditioning matrix ${\bm{P}}$ to be subject to sparsity constraints as well.
Algebraic preconditioners do not assume any additional problem information and compute the preconditioning matrix solely based on the structure and values of the matrix ${\bm{A}}$ in contrast to methods taking into account the underlying problem domain.
@@
In the simplest case, the Jacobi preconditioner approximates ${\bm{P}}$ using a diagonal matrix. More advanced methods such as incomplete factorizations or the Gauss-Seidel method compute approximate factorizations of the original system ${\bm{A}}$ which implicitly give rise to preconditioners as they allow efficient inversions (benzi2002preconditioning; pearson2020preconditioners). Most commonly, the approximation utilizes triangular factorizations since efficient inversion in at most $\mathcal{O}(n^{2})$ operations can be achieved and sparsity can be exploited (davis2016survey). However, It is also possible to approximate the matrix ${\bm{P}}^{-1}$ directly (benzi1999comparative).
@@ 2.2 Graph neural networks
Graph neural networks (GNNs) are a recently popularized family of neural network architectures designed to process data represented on unstructured grids or graphs (bronstein2021geometric). A graph is a tuple $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consisting of vertices (or nodes) $\mathcal{V}$ and a bivariate relation of edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. With slight abuse of notation, we refer to the edge features of the edge $e=(i,j)\in\mathcal{E}$ as ${\bm{e}}_{ij}\in\mathbb{R}^{m}$ and node features of node $i\in\mathcal{V}$ as ${\bm{n}}_{i}\in\mathbb{R}^{k}$. The message-passing framework updates the initial edge and node features – representing the input data – in each layer $l$ as follows:
${\bm{e}}_{ij}^{l+1}=\psi_{\bm{\theta}}({\bm{e}}_{ij}^{l},{\bm{n}}_{i}^{l},{\bm{n}}_{j}^{l})$ (2a)
$m_{i}^{l+1}=\bigoplus_{j\in\mathcal{N}(i)}{\bm{e}}_{ji}^{l+1}$ (2b)
${\bm{n}}_{i}^{l+1}=\phi_{\bm{\theta}}({\bm{n}}_{i}^{l},m_{i}^{l+1}).$ (2c)
@@ 3 Method
Throughout this paper, we are interested in solving many linear equation systems that share an underlying structure. Instead of having direct access to a probability distribution over problems, we obtain a dataset $D$ with a finite number of samples ${\bm{A}}_{1},{\bm{A}}_{2},\dots\bm{A}_{|D|}$. As previously stated, we assume each ${\bm{A}}_{i}$ to be square, non-singular and sparse.
@@ 3.1 Network architecture
The goal of this paper is to learn a mapping $f_{\bm{\theta}}$, parameterized by a graph neural network, that takes a non-singular matrix ${\bm{A}}\in\text{GL}(n,\mathbb{R})$ as input and outputs the corresponding preconditioner ${\bm{P}}$ for the equation system. In order to ensure that the preconditioning system is easily invertible – such that we can apply it within the GMRES algorithm – we learn the LU factorization of the preconditioner ${\bm{P}}$ instead, by mapping ${\bm{A}}$ to two real matrices ${\bm{L}}$ and ${\bm{U}}$. The output matrices are lower and upper triangular, respectively, with non-zero diagonal entries to ensure their invertibility as discussed later. For notational simplicity, we do not explicitly mention the dependence of the factors on the original matrix ${\bm{A}}$ and the neural network parameters ${\bm{\theta}}$ when the relationship is clear from the context.
Sparsity  Further, the two output factors, ${\bm{L}}$ and ${\bm{U}}$, are subject to sparsity constraints. Similar to incomplete LU factorization methods without fill-ins ILU(0), we enforce the same sparsity patterns to the elements of ${\bm{L}}$ and ${\bm{U}}$ as the original matrix ${\bm{A}}$. This constraint is directly encoded in the network architecture as we discuss next.
@@
We treat the system ${\bm{A}}$ as the adjacency matrix of a corresponding graph that we use directly within the message-passing framework. This transformation is known in the literature as Coates graph representation is depicted in Figure 1 (doob1984applications). Similar to hausner2023neural and tang2022graph, we use node features describing the structural properties of the corresponding rows and columns of the matrix. We use the non-zero elements of ${\bm{A}}$ as the corresponding edge features of the graph.
Compared to previous approaches which only output a single lower triangular factor, our model outputs two separate matrices ${\bm{L}}$ and ${\bm{U}}$. hausner2023neural add positional encodings about the output matrix by explicitly changing the Coates graph representation. This allows them to output the lower triangular factor as the message passing is only executed over this set of edges. In contrast, we add positional encodings to the input data by adding an additional edge feature to each directed edge in the graph. This indicates whether the final output embedding of the corresponding edge belongs to the upper- or lower-triangular part (±$\pm 1$ respectively) but does not change the underlying graph structure used for message passing.
@@
Invertibility  We require the predicted factors ${\bm{L}}$ and ${\bm{U}}$ of the preconditioner obtained from the forward pass through the model to be invertible in order to obtain a valid preconditioner for the GMRES method during inference. Therefore, we need to enforce non-zero elements on each diagonal.
@@
However, it is still possible that the output of the GNN results in a singular matrix as the final edge representations, obtained from message passing, can result in zero diagonal elements. Through additional measures, we avoid this in both factors and enforce non-singularity. Inspired by the classical LU factorization, we enforce unit diagonal elements for the ${\bm{U}}$ factor. For the diagonal elements in the lower triangular factor ${\bm{L}}$ we are using the activation function
$\zeta(x)=\begin{cases}x,&\text{if }|x|>\epsilon,\\ \epsilon,&\text{if }0\leq x\leq\epsilon,\\ -\epsilon,&\text{if }-\epsilon\leq x<0,\\ \end{cases}$ (3)
for a chosen small value $\epsilon=10^{-4}$. This guarantees that the output matrix ${\bm{L}}$ is invertible as diagonal elements with a small magnitude are shifted away from zero. However, this function is discontinuous at $x=0$ making it difficult to optimize using gradient-based methods. Therefore, we use the following continuous approximation of the activation function
$\hat{\zeta}(x)=x\cdot\left(1+\exp\left(-\left|\frac{4x}{\epsilon}\right|+2\right)\right)$ (4)

### v2

@@ Abstract
Incomplete LU factorizations of sparse matrices are widely used as preconditioners in Krylov subspace methods to speed up solving linear systems. Unfortunately, computing the preconditioner itself can be time-consuming and sensitive to hyper-parameters. Instead, we replace the hand-engineered algorithm with a graph neural network that is trained to approximate the matrix factorization directly. To apply the output of the neural network as a preconditioner, we propose an output activation function that guarantees that the predicted factorization is invertible. Further, applying a graph neural network architecture allows us to ensure that the output itself is sparse which is desirable from a computational standpoint. We theoretically analyze and empirically evaluate different loss functions to train the learned preconditioners and show their effectiveness in decreasing the number of GMRES iterations and improving the spectral properties on synthetic data. The code is available at https://github.com/paulhausner/neural-incomplete-factorization.
@@ 1 Introduction
The generalized minimal residual method (GMRES) (saad86gmres) is one of the most popular iterative methods to solve large-scale and sparse linear equation systems of the form ${\bm{A}}{\bm{x}}={\bm{b}}$. Throughout the paper, we assume the square system matrix ${\bm{A}}$ to be real-valued and full rank. Therefore, the unique solution to the equation system is given by ${\bm{A}}^{-1}{\bm{b}}$. However, inverting the matrix directly, or equivalently computing a full matrix factorization, scales computationally poorly and suffers from numerical instabilities. Instead, iterative Krylov subspace methods, such as GMRES, which refine an approximation of the solution in each step are the most common solving technique for large-scale and sparse matrices. The convergence of these methods depends on the conditioning and singular value distribution of the system matrix and the right-hand side ${\bm{b}}$. Therefore, the choice of appropriate preconditioner – designed to improve the system properties – is critical to obtain a fast and accurate solution for the equation system 10saad.
@@
One of the most common choices of preconditioners is the incomplete LU factorization (ILU). As the name suggests, the matrix ${\bm{A}}$ is factorized into lower (${\bm{L}}$) and upper (${\bm{U}}$) triangular factors allowing an efficient matrix inversion. To decrease the storage and computational time not all elements of the factorization are computed for the preconditioner leading to a sparse but incomplete factorization. The computation of the preconditioner is time-consuming, difficult to parallelize, and can suffer from numerical instabilities (benzi2002preconditioning; scott2023algorithms). In this work, we replace the numerical computation of the incomplete LU factors with a graph neural network (GNN). The GNN learns to predict the corresponding incomplete factors of the matrix directly by training against data. The main contributions of our paper are the following:
  • We design a GNN architecture that outputs a sparse and non-singular incomplete LU factorization.
@@
  • We theoretically analyze existing loss functions from the literature and derive their connection to large and small singular values.
@@
  • Building on these insights, we derive a novel loss function for training learned preconditioners that combines the benefits of prior approaches.
@@
In numerical experiments, we validate the effectiveness of our model in reducing GMRES iterations in combination with the different loss functions and validate the theoretical results on a synthetic dataset.
@@
Related work  Most similar to our work, hausner2023neural and li2023learning both learn an incomplete factorization for the conjugate gradient method. trifonov2024learning extend this to combine data-driven and classical algorithms by correcting the output of the incomplete Cholesky factorization with a learned component. However, all of these methods assume the matrix to be symmetric and positive definite and utilize the incomplete Cholesky factorization as the underlying preconditioning technique. In contrast, we rather focus on the more general incomplete LU factorization, which does not require positive definiteness or symmetry, within the GMRES algorithm similar to the recently proposed GraphPAN (yusuf2024constructing).
@@
For the class of sparse approximate inverse preconditioners, baankestad2024ising learn the sparsity pattern of the approximate inverse matrix for which the preconditioner can be computed efficiently. li2024generative instead propose to learn auto-encoder-based generative models to generate preconditioners. In this paper, we focus instead on learning factorized preconditioners that allow easy inversion rather than learning the inverse matrix directly.
@@ 2 Background
We start by providing a brief overview of the GMRES algorithm, emphasizing the critical role of preconditioners in accelerating convergence. Then, we describe graph neural networks which parameterize the mapping to learn data-driven preconditioners.
@@ 2.1 GMRES algorithm
We focus on GMRES (saad86gmres), a popular and widely adopted iterative Krylov subspace method to solve general linear equation systems of the form ${\bm{A}}{\bm{x}}={\bm{b}}$ where ${\bm{A}}$ is a $n\times n$ non-singular and real valued matrix, i.e. ${\bm{A}}\in$ GL$(n,\mathbb{R})$. The core idea of Krylov subspace methods is to iteratively refine an approximation of the solution to the problem starting from an initial guess ${\bm{x}}_{0}$. At each iteration, the approximate solution is computed by minimizing the Euclidean norm of the residual vector within the corresponding Krylov subspace:
@@
where the $k$-th Krylov subspace $\mathcal{K}_{k}({\bm{A}},{\bm{r}}_{0})=\text{span}\{{\bm{r}}_{0},{\bm{A}}{\bm{r}}_{0},{\bm{A}}^{2}{\bm{r}}_{0},\dots,{\bm{A}}^{k-1}{\bm{r}}_{0}\}$ of ${\bm{A}}$ is generated by the initial residual ${\bm{r}}_{0}={\bm{b}}-{\bm{A}}{\bm{x}}_{0}$ (10saad). By adding elements ${\bm{A}}^{i}{\bm{r}}_{0}$ to the subspace basis, the dimensionality of the Krylov subspace increases with each iteration. [^1] If the vector is linearly dependent to the basis the dimensionality remains the same but the exact solution to the linear equation system can be found within the existing subspace. This allows GMRES to further reduce the minimal residual norm in Equation (1) with more iterations. The GMRES algorithm is guaranteed to find the exact solution to the original linear equation system in at most $n$ steps.
@@
Given the previous solution ${\bm{x}}_{k}$ the next iterate can be computed efficiently by solving a reduced unconstrained system in the subspace. Each iteration consists of two main steps. In the first step of the algorithm, an orthonormal basis for the subspace $\mathcal{K}_{k}$ is computed. This can be achieved, for example, using Arnoldi’s algorithm which is based on the Gram-Schmidt procedure. In the second step, an unconstrained least squares problem is solved via the QR factorization of the subspace basis obtained in the first step. Based on these two steps, the solution of Equation (1) can be constructed efficiently. The detailed algorithm is shown in Appendix A.
@@
Preconditioning  Arguably, the most important design choice for any Krylov subspace method is the choice of preconditioner (benzi2002preconditioning). The goal of preconditioning is to replace the original linear equation system with a new preconditioned system that exhibits a better clustering of singular values, which in turn can lead to faster convergence of the iterative scheme (golub2013matrix). For a non-singular and easy-to-invert matrix ${\bm{P}}$, the linear equation system ${\bm{A}}{\bm{P}}^{-1}{\bm{y}}={\bm{b}}$ is solved instead of the original problem. The solution to the original problem is then given by ${\bm{x}}={\bm{P}}^{-1}{\bm{y}}$ [^2] Here, we use right preconditioning, other formulations use left- or split-preconditioning instead and often lead to similar but not equivalent results.
Constructing a preconditioner involves trading off the time required to compute the preconditioner itself and the resulting speedup in the iterative scheme (benzi2002preconditioning). Since the original system ${\bm{A}}$ is typically sparse, we restrict the preconditioning matrix ${\bm{P}}$ to be subject to sparsity constraints as well.
@@
Algebraic preconditioners do not assume any additional problem information and compute the preconditioning matrix solely based on the structure and values of the matrix ${\bm{A}}$. In contrast, geometric methods take into account the underlying problem domain. In the simplest case, the Jacobi preconditioner approximates the matrix ${\bm{A}}$ using a diagonal matrix. More advanced methods such as incomplete factorizations or the Gauss-Seidel method compute approximate factorizations of the original system ${\bm{A}}$ which implicitly give rise to preconditioners as they allow efficient inversions (benzi2002preconditioning; pearson2020preconditioners). Most commonly, the approximation utilizes triangular factorizations since efficient inversion in at most $\mathcal{O}(n^{2})$ operations can be achieved and sparsity can be exploited (davis2016survey). However, It is also possible to approximate the preconditioning matrix ${\bm{P}}^{-1}$ directly (benzi1999comparative).
@@ 2.2 Graph neural networks
Graph neural networks (GNNs) are a recently popularized family of neural network architectures designed to process data represented on unstructured grids or graphs (bronstein2021geometric).
A graph is a tuple $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consisting of vertices (or nodes) $\mathcal{V}$ and a bivariate relation of edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. With slight abuse of notation, we refer to the edge features of the edge $e=(i,j)\in\mathcal{E}$ as ${\bm{e}}_{ij}\in\mathbb{R}^{m}$ and node features of node $i\in\mathcal{V}$ as ${\bm{n}}_{i}\in\mathbb{R}^{k}$. The message-passing framework updates the initial edge and node features – representing the input data – in each layer $l$ as follows:
${\bm{e}}_{ij}^{l+1}=\psi_{\bm{\theta}}({\bm{e}}_{ij}^{l},{\bm{n}}_{i}^{l},{\bm{n}}_{j}^{l}),$ (2a)
${\bm{m}}_{i}^{l+1}=\bigoplus_{j\in\mathcal{N}(i)}{\bm{e}}_{ji}^{l+1},$ (2b)
${\bm{n}}_{i}^{l+1}=\phi_{\bm{\theta}}({\bm{n}}_{i}^{l},{\bm{m}}_{i}^{l+1}).$ (2c)
@@ 3 Method
Throughout this paper, we are interested in solving linear equation systems that share an underlying structure. Instead of having direct access to a probability distribution over the problems from a specific problem class, we obtain a dataset $\mathcal{D}$ with a finite number of i.i.d. samples ${\bm{A}}_{1},{\bm{A}}_{2},\dots\bm{A}_{|\mathcal{D}|}$. As previously stated, we assume each ${\bm{A}}_{i}$ to be square, non-singular and sparse.
@@ 3.1 Network architecture
The goal of this paper is to learn a mapping $f_{\bm{\theta}}$, parameterized by a graph neural network, that takes a non-singular matrix ${\bm{A}}\in\text{GL}(n,\mathbb{R})$ as input and outputs the corresponding preconditioner ${\bm{P}}$ for the equation system. In order to ensure that the preconditioning system is easily invertible – such that we can apply it within the GMRES algorithm – we learn the LU factorization of the preconditioner ${\bm{P}}$ instead, by mapping ${\bm{A}}$ to two real matrices ${\bm{L}}$ and ${\bm{U}}$. The output matrices are lower and upper triangular, respectively, with non-zero diagonal entries to guarantee their invertibility by utilizing a suitable activation function. For notational simplicity, we do not explicitly mention the dependence of the factors on the original matrix ${\bm{A}}$ and the neural network parameters ${\bm{\theta}}$ when the relationship is clear from the context.
Sparsity  The two output factors, ${\bm{L}}$ and ${\bm{U}}$, are subject to sparsity constraints. Similar to incomplete LU factorization methods without fill-ins, ILU(0), we enforce the same sparsity patterns to the elements of ${\bm{L}}$ and ${\bm{U}}$ as the original matrix ${\bm{A}}$. This sparsity constraint is directly encoded in the graph neural network architecture.
@@
We treat the system ${\bm{A}}$ as the adjacency matrix of a corresponding graph that we use directly within the message-passing framework. This transformation is known in the literature as Coates graph representation is depicted in Figure 1 (doob1984applications). Similar to hausner2023neural and tang2022graph, we use node features describing the structural properties of the corresponding rows and columns of the matrix. We use the non-zero elements of ${\bm{A}}$ as the input edge features of the graph.
Compared to previous approaches which only output a single lower triangular factor, our model outputs two separate matrices ${\bm{L}}$ and ${\bm{U}}$. hausner2023neural incorporate positional encoding by modifying the Coates graph representation, allowing them to output the lower triangular factor as the message passing is only executed over this subset of edges. In contrast, we add positional encoding directly to the input data by adding an edge feature to each directed edge in the graph. This feature indicates whether the final output embedding of the corresponding edge belongs to the upper- or lower-triangular part (±$\pm 1$ respectively) without changing the underlying graph structure used for message passing.
@@
Invertibility  To obtain a valid preconditioner for the GMRES method during inference, the predicted factors ${\bm{L}}$ and ${\bm{U}}$ from the model’s forward pass must be invertible. Therefore, we need to enforce non-zero elements on each diagonal.
@@
However, it is still possible that the output of the GNN results in a singular matrix as the final edge representations, obtained from message passing, can result in zero diagonal elements. Through additional measures, we avoid this in both factors and enforce non-singularity. Inspired by the classical LU factorization, we enforce a unit diagonal for the ${\bm{U}}$ factor. For the diagonal elements in the lower triangular factor ${\bm{L}}$ we are using the activation function
$\zeta_{\epsilon}(x)=\begin{cases}x,&\text{if }|x|>\epsilon,\\ \epsilon,&\text{if }0\leq x\leq\epsilon,\\ -\epsilon,&\text{if }-\epsilon\leq x<0.\\ \end{cases}$ (3)
In our experiments, we use the fixed hyper-parameter $\epsilon=10^{-4}$. This guarantees that the output matrix ${\bm{L}}$ is invertible as diagonal elements with a small magnitude are shifted away from zero. However, this function is discontinuous at $x=0$ making it difficult to optimize using gradient-based methods. Therefore, we use the following continuous approximation of the activation function:
$\hat{\zeta}_{\epsilon}(x)=x\cdot\left(1+\exp\left(-\left|\frac{4x}{\epsilon}\right|+2\right)\right)$ (4)